OpenStudy (anonymous):

Mr. Whipple wants to blend two teas, regular and all-spice, whose wholesale cost is $0.90/lb and $1.20/lb respectively. He wil sell the mixture at $1.65/lb and wishes to make a 50% profit over wholesale cost. What should the ratio of regular to all-spice be to accomplish this?

OpenStudy (anonymous):

hey guys

OpenStudy (anonymous):

i need help

OpenStudy (mathmale):

Hey, goel, what have you been able to do so far towards solving this problem?

OpenStudy (anonymous):

.... nothing im in an advanced class and dont know to do this

OpenStudy (mathmale):

This problem is a bit harder than those I'm used to in that we don't know the constraints on x and y, other that x and y must both be positive. Generally, we're told that we need to make P pounds of mixture selling at $1.65 per pound, which is a constraint: x + y = P. Here we don't have the luxury of having that info, at least not explicitly. Does this ring a bell at all? We need to figure out how to include a 50% profit into our calculations.

OpenStudy (mathmale):

In general, profit = Total selling price - (Total purchase price and costs).

OpenStudy (anonymous):

im confused

OpenStudy (mathmale):

"I'm confused" is a natural response. But I'm hoping you can turn your confusion into some good questions that might lead us to a solution. What do YOU think we need to know to solve this problem? I've already given you one hint: the selling price needs to be 50% higher than the price the seller pays wholesale for these two teas.

OpenStudy (mathmale):

@cgoel: we do need to choose at least 2 different variables to represent the unknowns here. What unknowns are we discussing here?

OpenStudy (mathmale):

Let x = the number of (?) of ( ? ) Let y = the number of (?) of ( ? )

OpenStudy (mertsj):

1.65(x+y)=1.5(.9x+1.2y) 1.65x+1.65y=1.35x+1.8y .30x=.15y x/y=.15/.30 x/y=1/2

OpenStudy (mathmale):

@cgoel: Mertsj has done a great job here of setting up a relationship between x and y. His 1.5 includes the 50% markup on wholesale. His (x+y) denotes the total number of pounds of the tea mixture. He recognizes that we do not need to find x and y, but only their ratio. His result, x/y = 1/2, tells us that the number of pounds of the cheaper tea should be 1/2 of the number of pounds of the more expensive tea. I ask that you read through this and ask questions about whatever remains unclear to you.

OpenStudy (anonymous):

thank you guys!

OpenStudy (mathmale):

"What should the ratio of regular to all-spice be to accomplish this?"

OpenStudy (mathmale):

"Ratio" is the key word here.